OpenStudy (anonymous):

Find the equation of the circle with center at (3, 2) and through the point (5, 4).

OpenStudy (anonymous):

@Kainui @Nurali

OpenStudy (bibby):

What is the equation of a circle? What is the radius of this circle?

OpenStudy (anonymous):

Isn't it Pythagorean?

OpenStudy (anonymous):

And I have no clue what the radius is

OpenStudy (bibby):

The radius is the distance from the center to any point on the circle. In this case, one of the points on the circle is (5,4) I think the distance formula would work

OpenStudy (anonymous):

Im not sure if the question is asking for the equation to solve it or if its asking for the answer through using the distance formula. Can you tell which its asking for?

OpenStudy (bibby):

It's asking for an equation. First, find the equation for a circle

OpenStudy (anonymous):

I'd want to use this correct? (x – h)2 + (y – k)2 = r2

OpenStudy (bibby):

Yes. What is h, k?

OpenStudy (bibby):

I meant what are the numbers that fit h and k lol

OpenStudy (anonymous):

Oh sorry. 3 and 2 (h and k)

OpenStudy (bibby):

So now we have the following:\[\huge {(x – 3)}^2 + {(y – 2)}^2 = r^2\]

OpenStudy (bibby):

we want to find the radius. now recall that the distance from 5,4 to 3,2 is the radius

OpenStudy (anonymous):

So we need to use the distance formula then?

OpenStudy (bibby):

Yes.

OpenStudy (anonymous):

I got 2.82 (5-3)^2+(4-2)^2 2^2+2^2 = 8 sqr. root of 8 being 2.82

OpenStudy (bibby):

You can leave it as sqrt(8) because look at what happens when we plug it in\[\huge {(x – 3)}^2 + {(y – 2)}^2 = \sqrt{8}^2\]\[\huge {(x – 3)}^2 + {(y – 2)}^2 = \sqrt{8}^ \cancel 2\]\[\huge {(x – 3)}^2 + {(y – 2)}^2 =8\]

OpenStudy (anonymous):

Oh alright so then our answer would be the last one then?

OpenStudy (bibby):

yup

OpenStudy (anonymous):

Thank you for your help. I might have another one soon if you're willing to help.

OpenStudy (bibby):

Sure, close this question and open a new one.